part a A 12.0 L gas cylinder is filled with 7.80 moles of gas. The tank...

part a

A 12.0 L gas cylinder is filled with 7.80 moles of gas. The tank is stored at 51 ∘C . What is the pressure in the tank? Express your answer to three significant figures and include the appropriate units.

part b

A 300. L kiln is used for vitrifying ceramics. It is currently operating at 1195 ∘C , and the pressure is 0.9000 atm . How many moles of air molecules are within the confines of the kiln?

Express your answer to three significant figures and include the appropriate units.

part c

A 18.0 L gas cylinder has been filled with 5.80 moles of gas. You measure the pressure to be 3.30 atm . What is the temperature inside the tank?

Express your answer to three significant figures and include the appropriate units.
Answer #1

A)

Given:

V = 12.0 L

n = 7.80 mol

T = 51.0 oC

= (51.0+273) K

= 324 K

use:

P * V = n*R*T

P * 12 L = 7.8 mol* 0.08206 atm.L/mol.K * 324 K

P = 17.3 atm

Answer: 17.3 atm

B)

Given:

P = 0.9 atm

V = 300.0 L

T = 1195.0 oC

= (1195.0+273) K

= 1468 K

find number of moles using:

P * V = n*R*T

0.9 atm * 300 L = n * 0.08206 atm.L/mol.K * 1468 K

n = 2.24 mol

Answer: 2.24 mol

c)

Given:

P = 3.3 atm

V = 18.0 L

n = 5.80 mol

use:

P * V = n*R*T

3.3 atm * 18 L = 5.8 mol* 0.08206 atm.L/mol.K * T

T = 125 K

Answer: 125 K